# Data Augmentation for Traffic Classification

### 19 Jan 2024 v1

@@ Abstract
Data Augmentation (DA)—enriching training data by adding synthetic samples—is a technique widely adopted in Computer Vision (CV) and Natural Language Processing (NLP) tasks to improve models performance. Yet, DA has struggled to gain traction in networking contexts, particularly in Traffic Classification (TC) tasks. In this work, we fulfill this gap by benchmarking $18$ augmentation functions applied to $3$ TC datasets using packet time series as input representation and considering a variety of training conditions. Our results show that ($i$) DA can reap benefits previously unexplored with ($ii$) augmentations acting on time series sequence order and masking being a better suit for TC and ($iii$) simple latent space analysis can provide hints about why augmentations have positive or negative effects.
@@ 1 Introduction
In this paper, we fill these gaps by providing a comprehensive evaluation of “hand-crafted” augmentations—transformations designed based on domain knowledge—applied to packets time series typically used as input in TC. Given the broad design space, we defined research goals across multiple dimensions. First of all, we selected a large pool of 18 augmentations across 3 families (amplitude, masking, and sequence) which we benchmark both when used in isolation as well as when multiple augmentations are combined (e.g., via stacking or ensembling). Augmentations are combined with original training data via different policies (e.g., replacing training data with augmentation, adding augmented data to each training step, or pre-augmenting the dataset before training). We also included scenarios where imbalanced datasets are re-balanced during training to give more importance to minority classes. Last, we dissected augmentations performance by exploring their geometry in the classifiers latent space to pinpoint root causes driving performance. Our experimental campaigns were carried over 2 mid-sized public datasets, namely $\mathtt{MIRAGE}$-$19$ and $\mathtt{MIRAGE}$-$22$ (up to 20 classes, 64k flows), and a larger private dataset (100 classes, 2.9M flows). We summarize our major findings as follows:
@@ 2.1 Data augmentation in traditional machine learning tasks
Supervised modeling and DA. In a nutshell, a supervised model is a function $\varphi:{\bf x}\in\mathcal{X}\rightarrow y\in\mathcal{Y}$ mapping an input $\bf x$ to its label $y$. Training such models corresponds to discovering a good function $\varphi(\cdot)$ based on a training set. When performing DA, the training set is enlarged by new samples ${\bf x^{\prime}}=\textrm{Aug}({\bf x})$ created by altering original samples ${\bf x}$—these transformations act directly in the input space $\mathcal{X}$ and the additional synthetic points contribute in defining $\varphi(\cdot)$ as much as the original samples. It follows that having a comprehensive understanding of samples/classes properties and their contribution to models training is beneficial for designing effective augmentations, i.e., transformations enabling higher classification performance.
@@
In traditional ML, a popular transformation technique is Synthetic Minority Over-sampling TEchnique (SMOTE) [7]. This approach generates new samples by interpolating the nearest neighbors of a given training sample. To address class imbalance, SMOTE is often employed with a sampling mechanism that prioritizes minority classes [16, 17].
In CV, several image transformations have been proposed to improve samples variety while preserving classes semantics. These transformations operate on colors (e.g., contrast and brightness changes, gray scaling) and geometry (e.g., rotation, flipping, and zooming), or via filters (e.g., blurring with Gaussian kernel) and masks (e.g., randomly set to zero a patch of pixels). Furthermore, transformations like CutMix [52] and Mixup [53] not only increase samples variety but also classes variety by creating synthetic classes from a linear combination of existing ones. The rationale behind this approach is that introducing new artificial classes sharing similarities with the true classes, the classification task becomes intentionally more complex, thereby pushing the training process to extract better data representations. Empirical validations of DA techniques in CV have consistently demonstrated their effectiveness across a diverse range of datasets, tasks, and training paradigms [8, 18, 32]. As a result, DA has become a ubiquitous component in the CV models training pipelines.
@@ 2.2 Data augmentation in traffic classification
TC is commonly operated using either packet time series (e.g., size, direction, Inter Arrival Time (IAT), etc., of the first 10-30 packets of a flow) or payload bytes (e.g., the first 784 bytes of a flow, possibly gathered by concatenating payload across different packets) arranged as 2d matrices. Recent literature also considers combining both input types into multi-modal architectures [2, 4, 24].
Such input representation and datasets exhibit three notable distinctions when compared to data from other ML/DL disciplines. First, TC datasets show significant class imbalance—this is a “native” property of network traffic as different applications enjoy different popularity and traffic dynamics while, for instance, many CV datasets are balanced. Second, TC input representation is typically “small” to adhere to desirable system design properties—network traffic should be ($i$) early classified i.e., the application associated to a flow should be identified within the first few packets of a flow, and ($ii$) computational/memory resources required to represent a flow should be minimal as an in-network TC system needs to cope with hundreds of thousands for flow per second. Last, TC input data has weak semantics—the underlying application protocols (which may or not be known a priori) may not be easily interpreted from packet time series and visually inspecting those patterns is challenging even for domain experts.
@@
The combination of the above observations leads to have only a handful of studies adopting DA in TC. Rezaei et al. [33] created synthetic input samples by means of three hand-crafted DA strategies based on sampling multiple short sequences across the duration of a complete flow. Horowicz et al. [20] instead focused on a flowpic input representation—a 2d summary of the evolution of packets size throughout the duration of a flow—augmented by first altering the time series collected from the first 15s of a flow and used to compute the flowpic. While both studies show the benefit of DA, these strategies violate the early classification principle as they both consider multiple seconds of traffic, thus they are better suited for post-mortem analysis only. Conversely, Xie et al. [47] recently proposed some packet series hand-crafted DA to tackle data shifts arising when applying a model on network traffic gathered from networks different from the ones used to collect the training dataset. Specifically, inspired by TCP protocol dynamics, authors proposed five packets time series augmentations (e.g., to mimic a packet lost/retransmission one can replicate a value at a later position in the time series) showing that they help to mitigate data shifts. Yet, differently from [20, 33], the study in [47] lacks an ablation of each individual augmentation’s performance.
@@
Last, [42, 44, 49] investigate augmentations based on GAN methods when using payload bytes as input for intrusion detection scenarios, i.e., a very special case of TC where the classification task is binary. More recently instead, [22] compared GAN and diffusion model for generating raw payload bytes traces while [38, 39] instead leveraged GAN or diffusion models to generate 2D representations (namely GASF) of longer traffic flow signals for downstream traffic fingerprinting, anomaly detection, and TC.
@@ 2.3 Design space
As observed in TC literature [20, 33, 47], domain knowledge is key to designing efficient augmentations, yet ingenuity might not be enough as models are commonly used as “black boxes”, making it extremely challenging to establish a direct link between an augmentation technique and its impact on the final classification performance. For instance, rotation is considered a good image transformation as result of empirical studies. Likewise, generative models are trained on large image datasets but without an explicit connection to a classification task [34]—the design of the augmentation method itself is part of a trial-and-error approach and the definition of metrics quantifying the augmentation quality is still an open question.
One of the aspects to be considered to formulate such metrics is the variety introduced by the augmentations. Gontijo-Lopes et al. [11] propose metrics quantifying the distribution shift and diversity introduced by DA contrasting models performance with and without augmentation. Other literature instead focuses on mechanisms that can help defining desirable properties for augmentations. For instance, from the feature learning literature, [36, 54] find that DA induces models to better learn rare/less popular but good features by altering their importance, thus improving model generalization performance. Samyak et al. [21] find that optimization trajectories are different when training on different augmentations and propose to aggregate the weights of models trained on different augmentations to obtain a more uniform distribution of feature patches, encouraging the learning of diverse and robust features.
@@
Self-supervision and contrastive learning are DL training strategies that take advantage of augmentations by design. In a nutshell, contrastive learning consists in a 2-steps training process. First, a feature extractor is trained in a self-supervised manner with a contrastive loss function that pulls together different augmented “views” of a given sample while distancing them from views of other samples. Then, a classifier head is trained on top of the learned representation in a supervised manner using a few labeled samples—the better the feature representation, the lower the number of labeled samples required for training the head. Empirical studies have demonstrated the robustness of the feature representations learned with contrastive learning [8] and a few recent studies considered contrastive learning also in TC [15, 20, 40, 47].
@@
When we consider the specific case of using generative models to augment training data, we face a major challenge—generative models are not designed to target a specific downstream task [38, 39, 44]. While studies like [29] integrated a classifier in GAN training in the pursuit of improving the reliability of a GAN, how to properly link and train a generative model to be sensitive toward a downstream classification task is still an open question even in CV literature.
@@ 3 Our goals and methodology
Drawing insights from the existing literature review in Sec. 2, we undertake an empirical campaign to better understand hand-crafted DA applied in the input space for a specific task, namely TC. In essence, our research goals encompass the following goals:
@@
Figure 1: Input sample shape and related notation.
@@ 3.1 Benchmarking hand-crafted DA (G1)
In this study, we considered a set $\mathcal{A}$ of 18 augmentation functions denoted by $Aug(\cdot)$. These functions can be categorized into 3 families: 5 amplitude transformations, which introduce different type of jittering to the feature values (Table 1); 2 masking transformations, which force certain feature values to zero (Table 2); and 11 sequence transformations, which modify the order of feature values (Table 3). It is important to note that, given a sample $\rm{\bf x}$, amplitude augmentations are solely applied to either packet size or IAT while packet direction is never altered since the latter is a binary feature and does not have amplitude i.e., it can be $-1$ or $1$. On the contrary masking and sequence augmentations are applied to all features in parallel (e.g., if a transformation requires to swap $t=1$ with $t=6$, all three features are swapped accordingly $x_{(i,1)}\leftrightarrow x_{(i,6)}$ for $\forall i\in\{0..D-1\}$). For each augmentation, the tables report a reference example annotating its parametrization (if any).
By adopting such a large pool of augmentations our empirical campaign offers several advantages. First, we are able to investigate a broader range of design possibilities compared to previous studies. Second, it enables us to contrast different families and assess if any of them is more prone to disrupt class semantics. Considering the latter, TC literature [20, 33, 47] predominantly investigate sequence transformations (typically acting only on packet timestamp) with only [47] also experimenting with masking and amplitude variation, yet targeting scenarios where models are exposed to data shifts due to maximum segment size (MSS) changes, i.e., the network properties related to the training set are different from the ones of the test set.
@@
As described in Tables 1-3, each augmentation has some predefined static parameters [^1] These parameters are tuned via preliminary investigations. while the magnitude $\alpha$ is the single hyper-parameter controlling random sampling mechanisms contributing to defining the final transformed samples. To quantify augmentation sensitivity to $\alpha$, we contrast two scenarios following CV literature practice: a static value of $\alpha=0.5$ and a uniformly sampled value $\alpha\sim U[0,1]$ extracted for each augmented sample.
@@ 3.2 Training batches composition (G2)
In order to mitigate any undesirable shifts introduced by artificial samples, it is necessary to balance original and augmented samples. Yet, the way original and augmented samples are combined to form the augmented training set is a design choice. For instance, in TC literature [20, 33] augment the data before starting the training, while [47] augments mini-batches during the training process which is closer to what is done in the CV literature. In this work, we apply augmentation samples to a training mini-batch of size $B$, with the two policies sketched in Fig. 2. Replace substitutes an original sample $\rm{\bf x_{i}}$ with its augmentation by sampling from a Bernoulli(P=$P_{replace}$) random variable—during one training epoch, approximately a $P_{replace}$ fraction of the original data is “hidden”. Instead, Inject increases the batch size by augmenting each sample $N_{inject}$ times (e.g., in Fig. 2 the original batch size is doubled by setting $N_{inject}=1$).
@@ 3.3 Latent space geometry (G3)
Augmentation-vs-Test. ML methods operate on the assumption that training data serves as a “proxy” for test samples, i.e., the patterns learned on training data “generalize” to testing data as the two sets of data resemble each other properties. In this context, augmentations can be considered as a means of further push such data generalization by incorporating samples that resemble even more testing data compared to what is available in training data. However, it is important to empirically quantify this “effect” by measuring, for instance, the distance between augmented and test samples. In other words, we aim to quantify up to which extent augmented samples are better at mimicking test samples compared to the original data.
Augmentation-vs-Train. The performance of a feature extractor greatly depends on how well the feature extractor separates different classes in the latent space. Data augmentations play a role in shaping intra/inter-class relationships created by the feature extractor in the latent space. For instance, an augmentation that generates samples far away from the region of a class can disrupt class semantics—the augmentation is introducing a new behavior/mode making it hard for the classifier to be effective. At the same time, however, expanding the region of a class can be a beneficial design choice—augmentations that enable a better definition of class boundaries simplify the task of the classifier. Understanding such dynamics requires empirical observations, for instance, comparing the distance between original training data and augmented data. In other words, we aim to verify if augmentations yielding good performance are in a “sweet spot”: they create samples that are neither too close (i.e., introduce too little variety) nor too far (i.e., disrupt class semantics) from original samples.

### 23 Jan 2024 v2

@@ Abstract
Data Augmentation (DA)—enriching training data by adding synthetic samples—is a technique widely adopted in Computer Vision (CV) and Natural Language Processing (NLP) tasks to improve models performance. Yet, DA has struggled to gain traction in networking contexts, particularly in Traffic Classification (TC) tasks. In this work, we fulfill this gap by benchmarking $18$ augmentation functions applied to $3$ TC datasets using packet time series as input representation and considering a variety of training conditions. Our results show that ($i$) DA can reap benefits previously unexplored, ($ii$) augmentations acting on time series sequence order and masking are better suited for TC than amplitude augmentations and ($iii$) basic models latent space analysis can help understanding the positive/negative effects of augmentations on classification performance.
@@ 1 Introduction
In this paper, we fill these gaps by providing a comprehensive evaluation of “hand-crafted” augmentations—transformations designed based on domain knowledge—applied to packets time series typically used as input in TC. Given the broad design space, we defined research goals across multiple dimensions. First of all, we selected a large pool of 18 augmentations across 3 families (amplitude, masking, and sequence) which we benchmark both when used in isolation as well as when multiple augmentations are combined (e.g., via stacking or ensembling). Augmentations are combined with original training data via different batching policies (e.g., replacing training data with augmentation, adding augmented data to each training step, or pre-augmenting the dataset before training). We also included scenarios where imbalanced datasets are re-balanced during training to give more importance to minority classes. Last, we dissected augmentations performance by exploring their geometry in the classifiers latent space to pinpoint root causes driving performance. Our experimental campaigns were carried over 2 mid-sized public datasets, namely $\mathtt{MIRAGE}$-$19$ and $\mathtt{MIRAGE}$-$22$ (up to 20 classes, 64k flows), and a larger private dataset (100 classes, 2.9M flows). We summarize our major findings as follows:
@@ 2.1 Data augmentation in traditional machine learning tasks
Supervised modeling and DA. In a nutshell, a supervised model is a function $\varphi:{\bf x}\in\mathcal{X}\rightarrow y\in\mathcal{Y}$ mapping an input $\bf x$ to its label $y$. Training such models corresponds to discover a good function $\varphi(\cdot)$ based on a training set. When performing DA, the training set is enlarged by adding new samples ${\bf x^{\prime}}=\textrm{Aug}({\bf x})$ created by altering original samples ${\bf x}$—these transformations act directly in the input space $\mathcal{X}$ and the additional synthetic samples contribute in defining $\varphi(\cdot)$ as much as the original ones. It follows that having a comprehensive understanding of samples/classes properties and their contribution to models training is beneficial for designing effective augmentations, i.e., transformations enabling higher classification performance.
@@
In traditional ML, Synthetic Minority Over-sampling TEchnique (SMOTE) [7] is a popular augmentation technique. This approach generates new samples by interpolating the nearest neighbors of a given training sample. To address class imbalance, SMOTE is often employed with a sampling mechanism that prioritizes minority classes [16, 17].
In CV, several image transformations have been proposed to improve samples variety while preserving classes semantics. These transformations operate on colors (e.g., contrast and brightness changes, gray scaling) and geometry (e.g., rotation, flipping, and zooming), or via filters (e.g., blurring with Gaussian kernel) and masks (e.g., randomly set to zero a patch of pixels). Furthermore, transformations like CutMix [51] and Mixup [52] not only increase samples variety but also increase classes variety by creating synthetic classes from a linear combination of existing ones. The rationale behind this approach is that by introducing new artificial classes sharing similarities with the true classes the classification task becomes intentionally more complex, thereby pushing the training process to extract better data representations. Empirical validations of DA techniques in CV have consistently demonstrated their effectiveness across a diverse range of datasets, tasks, and training paradigms [8, 18, 31]. As a result, DA has become a ubiquitous component in the CV models training pipelines.
@@ 2.2 Data augmentation in traffic classification
TC tasks usually rely on either packet time series (e.g., packet size, direction, Inter Arrival Time (IAT), etc., of the first 10-30 packets of a flow) or payload bytes (e.g., the first 784 bytes of a flow, possibly gathered by concatenating payload across different packets) arranged as 2d matrices. Recent literature also considers combining both input types into multi-modal architectures [2, 4, 24].
Such input representations and datasets exhibit three notable distinctions when compared to data from other ML/DL disciplines. First, TC datasets show significant class imbalance—this is a “native” property of network traffic as different applications enjoy different popularity and traffic dynamics while, for instance, many CV datasets are balanced. Second, TC input representation is typically “small” to adhere to desirable system design properties—network traffic should be ($i$) early classified, i.e., the application associated to a flow should be identified within the first few packets of a flow, and ($ii$) computational/memory resources required to represent a flow should be minimal as an in-network TC systems need to cope with hundreds of thousands of flow per second. Last, TC input data has weak semantics—the underlying application protocols (which may or not be known a priori) may not be easy to interpret even for domain experts when visually inspecting packet time series.
@@
The combination of the above observations leads to have only a handful of studies adopting DA in TC. Rezaei et al. [32] created synthetic input samples by means of three hand-crafted DA strategies based on sampling multiple short sequences across the duration of a complete flow. Horowicz et al. [20] instead focused on a flowpic input representation—a 2d summary of the evolution of packets size throughout the duration of a flow—augmented by first altering the time series collected from the first 15s of a flow and used to compute the flowpic. While both studies show the benefit of DA, these strategies violate the early classification principle as they both consider multiple seconds of traffic, thus they are better suited for post-mortem analysis only. Conversely, Xie et al. [46] recently proposed some packet series hand-crafted DA to tackle data shifts arising when applying a model on network traffic gathered from networks different from the ones used to collect the training dataset. Specifically, inspired by TCP protocol dynamics, authors proposed five packets time series augmentations (e.g., to mimic a packet lost/retransmission one can replicate a value at a later position in the time series) showing that they help to mitigate data shifts. Yet, differently from [20, 32], the study in [46] lacks from an ablation of each individual augmentation’s performance.
@@
Last, [41, 43, 48] investigate augmentations based on GAN methods when using payload bytes as input for intrusion detection scenarios, i.e., a very special case of TC where the classification task is binary. More recently, [22] compared GAN and diffusion model for generating raw payload bytes traces while [37, 38] instead leveraged GAN or diffusion models to generate 2D representations (namely GASF) of longer traffic flow signals for downstream traffic fingerprinting, anomaly detection, and TC.
@@ 2.3 Design space
As observed in TC literature [20, 32, 46], domain knowledge is key to design efficient augmentations. Yet, ingenuity might not be enough as models are commonly used as “black boxes”, making it extremely challenging to establish a direct link between an augmentation technique and its impact on the final classification performance. For instance, rotation is considered a good image transformation as result of empirical studies. Likewise, generative models are trained on large image datasets but without an explicit connection to a classification task [33]—the design of the augmentation method itself is part of a trial-and-error approach and the definition of metrics quantifying the augmentation quality is still an open question.
One of the aspects to be considered when formulating such metrics is the variety introduced by the augmentations. Gontijo-Lopes et al. [11] propose metrics quantifying the distribution shift and diversity introduced by DA contrasting models performance with and without augmentations. Other literature instead focuses on mechanisms that can help defining desirable properties for augmentations. For instance, from the feature learning literature, [35, 53] find that DA induces models to better learn rare/less popular but good features by altering their importance, thus improving model generalization performance. Samyak et al. [21] find that optimization trajectories are different when training on different augmentations and propose to aggregate the weights of models trained on different augmentations to obtain a more uniform distribution of feature patches, encouraging the learning of diverse and robust features.
@@
Self-supervision and contrastive learning are DL training strategies that take advantage of augmentations by design. In a nutshell, contrastive learning consists of a 2-steps training process. First, a feature extractor is trained in a self-supervised manner with a contrastive loss function that pulls together different augmented “views” of a given sample while distancing them from views of other samples. Then, a classifier head is trained on top of the learned representation in a supervised manner using a few labeled samples—the better the feature representation, the lower the number of labeled samples required for training the head. Empirical studies have demonstrated the robustness of the feature representations learned with contrastive learning [8] and a few recent studies investigated contrastive learning also in TC [15, 20, 39, 46].
@@
When we consider the specific case of using generative models to augment training data, we face a major challenge—generative models are not designed to target a specific downstream task [37, 38, 43]. While studies like [28] integrated a classifier in GAN training in the pursuit of improving the reliability the model, how to properly link and train a generative model to be sensitive to a downstream classification task is still an open question even in CV literature.
@@ 3 Our goals and methodology
Drawing insights from the literature reviewed in Sec. 2, we undertake a set of empirical campaigns to better understand hand-crafted DA when applied in the input space for TC task and address the following research goals:
@@
Figure 1: Input sample ${\bf x}$ shape and related notation.
@@ 3.1 Benchmarking hand-crafted DA (G1)
In this study, we considered a set $\mathcal{A}$ of 18 augmentation functions. These functions can be categorized into 3 families: 5 amplitude transformations, which introduce different type of jittering to the feature values (Table 1); 2 masking transformations, which force certain feature values to zero (Table 2); and 11 sequence transformations, which modify the order of feature values (Table 3). It is important to note that, given a sample $\rm{\bf x}$, amplitude augmentations are solely applied to either packet size or IAT while packet direction is never altered since the latter is a binary feature and does not have amplitude (i.e., it can be $-1$ or $1$). On the contrary, masking and sequence augmentations are applied to all features in parallel (e.g., if a transformation requires to swap $t=1$ with $t=6$, all features are swapped accordingly $x_{(i,1)}\leftrightarrow x_{(i,6)}$ for $\forall i\in\{0..D-1\}$). For each augmentation, Tables 1-3 report a reference example annotating its parametrization (if any).
By adopting such a large pool of augmentations our empirical campaign offers several advantages. First, we are able to investigate a broader range of design possibilities compared to previous studies. Second, it enables us to contrast different families and assess if any of them is more prone to disrupt class semantics. Considering the latter, TC literature [20, 32, 46] predominantly investigate sequence transformations (typically acting only on packet timestamp) with only [46] experimenting with masking and amplitude variation, yet targeting scenarios where models are exposed to data shifts due to maximum segment size (MSS) changes, i.e., the network properties related to the training set are different from the ones of the test set.
@@
As described in Tables 1-3, each augmentation has some predefined static parameters [^1] These parameters are tuned via preliminary investigations. while the magnitude $\alpha$ is the single hyper-parameter controlling random sampling mechanisms contributing to defining the final transformed samples. To quantify augmentations sensitivity to $\alpha$, we contrast two scenarios following CV literature practice: a static value of $\alpha=0.5$ and a uniformly sampled value $\alpha\sim U[0,1]$ extracted for each augmented sample.
@@ 3.2 Training batches composition (G2)
In order to mitigate any undesirable shifts introduced by artificial samples, it is necessary to balance original and augmented samples. Yet, the way original and augmented samples are combined to form the augmented training set is a design choice. For instance, in TC literature, [20, 32] augment the data before starting the training, while [46] augments mini-batches during the training process. In this work, we apply augmentation samples to a training mini-batch of size $B$, with the two policies sketched in Fig. 2. Replace substitutes an original sample $\rm{\bf x_{i}}$ with its augmentation by sampling from a Bernoulli(P=$P_{replace}$) random variable—during one training epoch, approximately a $P_{replace}$ fraction of the original data is “hidden”. Instead, Inject increases the batch size by augmenting each sample $N_{inject}$ times (e.g., in Fig. 2 the original batch size is doubled by setting $N_{inject}=1$).
@@ 3.3 Latent space geometry (G3)
Augmentation-vs-Test. ML methods operate on the assumption that training data serves as a “proxy” for test samples, i.e., the patterns learned on training data “generalize” to testing data as the two sets of data resemble each other properties. In this context, augmentations can be considered as a means for fostering data generalization by incorporating samples that resemble even more testing data compared to what is available in training data. However, it is important to empirically quantify this effect by measuring, for instance, the distance between augmented and test samples. In other words, we aim to quantify up to which extent augmented samples are better at mimicking test samples compared to the original data.
Augmentation-vs-Train. The performance of a feature extractor greatly depends on how well the feature extractor separates different classes in the latent space. Data augmentations play a role in shaping intra/inter-class relationships created by the feature extractor in the latent space. For instance, an augmentation that generates samples far away from the region of a class can disrupt class semantics—the augmentation is introducing a new behavior/mode making it hard for the classifier to be effective. At the same time, however, expanding the region of a class can be a beneficial design choice—augmentations that enable a better definition of class boundaries simplify the task of the classifier. Understanding such dynamics requires empirical observations, for instance, by comparing the distance between original training data and augmented data. In other words, we aim to verify if augmentations yielding good performance are in a “sweet spot”: they create samples that are neither too close (i.e., introduce too little variety) nor too far (i.e., disrupt class semantics) from original samples.